40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 656     Entry time: Thu Jul 10 19:12:07 2008
Author: Alberto 
Type: Update 
Category: General 
Subject: abs cavity length measurement experiment 
Yesterday morning, when I started, I found the IFO beam on a different position and the beam spot at the AS port looked very deformed. The overlapping with the secondary beam was not good enough to observe the beats anymore. Restoring the alignments of the interferometers did not work because, as John found out later, some of the photodiodes had offsets and gain which made the restore script ineffective. After resetting the parameters, we had to align every mirror of the interferometers and save the configuration twice. The second times was because on the first time the alignment had been done with the illuminators on. To avoid that in the future, John wrote an alarm to warn about the status of the lights.

After that we fixed the IFO beam, I had to realign the optics in the table to match the secondary beam to the IFO beam. I got the two beam overlapping and, even though the NPRO spot looked distorted, I could observe again some signal of the beat. To do that it was also necessary to have all the interferometer mirrors aligned so that we had more power from the ifo beam although it also made the spot flash. Ideally, to avoid the flashing (which we would also impede the PLL to work) we should work with the interferometer locked. Since that doesn't seem actually possible, we should just keep one of the ITM aligned and improve the beam matching so that we can observe the beats even with less power.

Today I spent the day trying to improve the alignement of the optics to observe the beats with only the ITM aligned, resetting the alignment of both beams with the ireses, with the Farady and all the rest. It was a rather long and tiring process but I think I'm close to the target and maybe tomorrow.
ELOG V3.1.3-